7 kips Calculate the magnitude and direction of the resultant of the shown forces. kips 50 o kips 10° 20°
Answer #1

net horizontal force

Fx = - 9 cos 29 - 7 cos 70 + 5 cos 40

Fx = - 6.435 kips

net vertical force

Fy = 5 sin 40 + 7 sin 70 + 9 sin 20

Fy = 12.87 kips

=========

a)

net magnitude of force

F^2 = Fx^2 + Fy^2

F = 14.39 kips

========

b)

direction

phi = 180 - arctan ( 12.87 / 6.435)

phi = 116.565 (ccw from +x axis)

=======

Comment in case any doubt, will reply for sure.. Goodluck